OpenStudy (anonymous):

find the slope of the tangent to the graph 3xy - 2x + 3y^2 = 5 at the point (2,1)

OpenStudy (zzr0ck3r):

so take the derivative \((3xy)'-2x'+(3y^2)'=0 \\3xy'+3y-2+6yy'=0\\y'=\frac{-3y+2}{3x+6y}\)

OpenStudy (anonymous):

same

OpenStudy (zzr0ck3r):

plug in for x,y...

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

wait I plug the x and y they give me for -3y+2 / 3x+6y

OpenStudy (zzr0ck3r):

\(\frac{-3(1)+2}{3(2)+6(1)}=\frac{-1}{12}\)

OpenStudy (anonymous):

ok yes , that is what I did